OpenStudy (anonymous):

Does the difference of two squares identity apply only to algebraic variables? Can we not express any known integer as difference of 2 squares?

OpenStudy (anonymous):

do you mean with integer solutions?

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

lets try it

OpenStudy (anonymous):

\[17=17\times 1\] \[(m+n)(m-n)=17\] \[m+n=17,m-n=1\] \[m = 9,n=8\] \[9^2-8^2=17\] works with 17

OpenStudy (anonymous):

\[10=5\times 2\] \[m+n=5,m-n=2\] oops this won't work because now \[m=\frac{7}{2}\]

OpenStudy (anonymous):

maybe \[10=10\times 1\] \[m+n=10,m-n=1\] nope because now \[m=\frac{11}{2}\]

OpenStudy (anonymous):

now it is more or less clear right?

OpenStudy (anonymous):

So does it only work with prime numbers?

OpenStudy (anonymous):

lets try this one \[15=5\times 3\] \[m+n=5,m-n=3\] \[m=4,n=1\] this one works \[15=4^2-1^2\]

OpenStudy (anonymous):

So is there no general rule to see which numbers this rule works for?

OpenStudy (anonymous):

in fact we can do it two ways \[15=15\times 1\] \[m+n=15,m-n=1\] \[m=8,n=7\] \[15=8^2-7^2\]

OpenStudy (anonymous):

no i was just pointing out that it is not a matter of "priime" or not

OpenStudy (anonymous):

lets try 12 \[12=4\times 3\] \[m+n=4,m-n=3\] \[2m=7\]no \[12=12\times 1\] \[m+n=12,m-n=1\] \[2m=13\] no \[12=6\times 2\] \[m+n=6,m-n=2\] \[2m=8\] yes \[m=4\] \[m=2\] \[12=4^2-2^2\]